A.p=1,2,3,4........49
Find a number in a.p such that sum of numbers before it is equal to sum of numbers after it....
Please answer fast...

Answer by rothauserc(4718)

the sum of an arithmetic progression is
:
S(n) = n * (a(1) + a(n)) / 2
:
we are given the following
:
n * (n + 1) / 2 = (49 - n - 1) * (n + 2 + 49) / 2
:
multiply both sides of = by 2 and simplify
:
n^2 + n = (48 - n) * (n + 51)
:
n^2 + n = -n^2 - 3n + 2448
:
2n^2 + 4n - 2448 = 0
:
divide both sides of = by 2
:
n^2 + 2n - 1224 = 0
:
use quadratic formula
:
n = (-2 + square root(2^2 + 4 * 1224)) / 2 = 34
n = (-2 - square root(2^2 + 4 * 1224)) / 2 = -36
:
we reject the negative solution and accept n = 34
:
**************************************
p = 35
:
check our answer
:
First sum = 34 * (1 + 34) / 2 = 595
Second sum = 14 * (36 + 49) / 2 = 595
:
our answer checks
*************************************
: